a 200m train starts from rest with acceleration 4m/s2. Head lights of its engine is switched on at t=60s and its tail lights at t=120s. Find the distance between the two events for an observer standing on platform.

Grade:11

7 Answers

Adarsh
6 years ago
Dea
Adarsh
6 years ago
the distance between the two events for an observer standing on platform will be 27880 meter.
rishu ranjan
6 years ago
1. Time (T), velocity (C) and angular momentum (h) are chosen as fundamental quantities instead of mass, length and time. In terms of these, the dimensions of mass would be : 
akshaj
6 years ago
it may be 36000 metres becauseif acceleration is 4 m/s^2 when time is 60sthe distance between two platform will be 2.3 km...
Rama sai rohan
6 years ago
S1 ,S2 is distance covered in 60s,120sS1=0.5×4×3600 =7200+200(train length)S2=0.5×4×14400 28800+200S2-S1=21600m
Gitanjali Rout
5 years ago
he distance between the two events for an observer standing on platform will be 27880 meter. ...........................................................................................................
Gitanjali Rout
5 years ago
S1 ,S2 is distance covered in 60s,120sS1=0.5×4×3600 =7200+200(train length)S2=0.5×4×14400 28800+200S2-S1=21600mbhj ////////////////////////////////////////